OpenStudy (acespeedfighter):

Which of the values for x and y make the equation 3x + 4y + 6 = 26 true? (4 points) x = 4, y = 2 x = 5, y = 6 x = 4, y = 3 x = 3, y = 5 3. (07.01) Which of the following sets shows all the numbers from the set {0, 0.5, 1, 2.5, 3} that make the inequality 2a + 3 ≥ 8 true? (4 points) {2.5, 3} {1, 2.5, 3} {0, 0.5, 1} {0.5, 1}

pooja195 (pooja195):

For both of these plug in your options into the equations and see what makes a true statement.

OpenStudy (acespeedfighter):

hmm....

pooja195 (pooja195):

#1 3(4) + 4(2) + 6 = 26 Your first option is x = 4, y = 2 So take 4 and 2 and sub them in for the variables (be sure to put them in the right place) If its true you should get 26=26 if it's not you'll get any number=26 So on an so forth if those don't work try your 2nd option. #2 2a + 3 ≥ 8 First option {2.5, 3} Plug those in for a and see if they make a true statement 2(2.5)+ 3 ≥ 8 2(3) + 3 ≥ 8 If they do good if they don't try the 2nd option etc

OpenStudy (acespeedfighter):

1. C 2. im not to sure but i think its B or A

pooja195 (pooja195):

1) C: 3(4) + 4(3) + 6 = 26 12+12+6=26 30=26 C is incorrect. 2) Don't guess actually try.

OpenStudy (acespeedfighter):

2. i really Confusing ;-;

OpenStudy (acespeedfighter):

1. A 2. D

OpenStudy (oswaldmurphy):

Do you still need help? He had to go.